OpenStudy (anonymous):

Workout in substitution method 2a+4b=-24 a=20-2b. list as no solution or infinite let me know if I have the sets correct.

OpenStudy (anonymous):

I have -14,1 and no solution

OpenStudy (anonymous):

\(2a+4b=-24\) \(a=20-2b\) If we divide the first equation by 2 we have \ \(a+2b=-12\) and the second equation is \(a+2b=20\) so there is no solution..

OpenStudy (anonymous):

thanks